OpenStudy (king):

A fish in an aquarium approaches the left wall at a rate of 3 m/s and observes a fly approaching it at 8 m/s. If the refractive index of water is 4/3, find the actual velocity of the fly.Explain with steps..pls help!!!:(

OpenStudy (anonymous):

Is it 4 m/s?

OpenStudy (king):

actual velocity of fly=3.75m/s

OpenStudy (anonymous):

We know that, Speed of light in air / Speed of light in water = 4/3 Or, 3 * 10^8 / x = 4/3 Or, x = 2.25 * 10^8 Now, 2.25*10^8 = 3 So, 3*10^8 = 3/(2.25 * 10^8) * 3*10^8 = 9/2.25 = 4

OpenStudy (anonymous):

Not getting how it is 3.75 m/s

OpenStudy (king):

wait hw is 2.25*10^8=3?

OpenStudy (anonymous):

I just equated for the sake of unitary method, since fish is in water. It may be wrong.

OpenStudy (anonymous):